# How to Securely Shuffle? A survey about Secure Shufflers for privacy-preserving computations

### v1

@@ 3.3.2. Correctness
Next to anonymity, a good shuffler should also satisfy correctness. This property covers various security definitions present in existing security papers. On the one hand, secure shufflers based on MPC (e.g., RPM (Lu and Kate, 2023)) prove a standard security property called “correctness” (i.e., the MPC protocol outputs the expected output). On the other hand, recent anonymous communication systems (e.g., Blinder (Abraham et al., 2020)) often guarantee censorship resistance. This second security guarantee is essential, but it does not prevent all the threats covered by the “correctness.” Taking inspiration from MPC, we propose the following explicit definition for shuffling correctness:
@@ 3.3.3. Sender disruption resistance
Several ML works studied the resilience of distributed training protocols to malicious behaviors (Chen et al., 2017; Alistarh et al., 2018). These works conceive protocols that succeed even in the presence of malicious or faulty agents. Protocol disruption is also possible in secure shuffling and is critical when thousands or millions of senders are involved. Controlling the honesty of millions of senders is too complex, so a form of disruption resistance is necessary. We propose the following definition regarding the resistance to malicious senders:
@@
This property provides tolerance to drop-outs (i.e., the disconnections of senders before the protocol completion), which drew much attention in the secure aggregation literature (Bell et al., 2020; Sabater et al., 2021; Bell et al., 2022b). Secure aggregation protocols such as (Bell et al., 2020, 2022b) only tolerate a fixed number of drop-outs, while sender disruption-resistant shufflers tolerate an unbounded number of drop-outs. Sender disruption resistance is stronger than drop-out tolerance since it also covers the reception of ill-formed data.
@@
Finally, Definition 3.7 specifies that the disruption can only be caused by missing or ill-formed messages to exclude Distributed Denial-of-Service (DDoS) attacks (i.e., flooding a server with a massive amount of requests to make it unavailable). These attacks (or any other network-level attack) are not specific to shufflers and can be mitigated using network security techniques (Mirkovic and Reiher, 2004).
@@ 3.3.4. Server disruption resistance
This property is called “robustness” in MPC protocols (Abraham et al., 2020). One may see it as an extra property because the number of servers is small enough to maintain a blocklist of untrustworthy servers. As for sender disruption resistance, server disruption resistance brings resilience to drop-outs.
@@ 3.4.1. Intersection attacks
These attacks (Wolinsky et al., 2013; Danezis and Serjantov, 2005) rely on the statistical analysis that can be made from several shuffle outputs to de-anonymize messages. Indeed, if a data owner participates in $k$ shuffles, computing the intersection of the round results might leak some information about this data owner. These attacks against shufflers are specific to the use case and the data being shuffled. For example, the intersection attack is useless if the shuffled messages are secret shares, as in Ishai et al. (Ishai et al., 2006) protocol, but the intersection may de-anonymize the messages if we shuffle cleartext data.
Wolinsky et al. (Wolinsky et al., 2013) propose a generic Buddy system compliant with any anonymization system to mitigate these attacks. The general idea is to ensure a large set of possible senders for a given message stream, so the intersection of multiple shuffles cannot lead to a de-anonymization. No result is available regarding the scalability of this generic mitigation technique for applications to privacy-preserving computations.
@@ 3.4.2. Trickle and flooding attacks
Serjantov et al. (Serjantov et al., 2003) describe two general attacks (grouped under the term “blending attacks”): trickle attacks and flooding attacks.
@@
The trickle attack is only possible on a system insecure against censorship (i.e., with no correctness guarantee as defined in Section 3). The flooding attack is only possible if a malicious server colludes with data owners and can be mitigated using an authentication scheme, so a single malicious user cannot flood a batch alone. Some secure shufflers (e.g., mix nets) must use anonymous authentication systems (Chase et al., 2014) to mitigate these attacks.
@@ 3.4.3. Traffic analysis attacks
These attacks (Murdoch and Danezis, 2005; Sun et al., 2015; Arp et al., 2015) have been extensively documented for mix networks: a global adversary (e.g., a state agency) observes the packet transmission between the agents to trace the path of the messages. From the beginning, anonymization networks such as Tor were presented as possibly insecure against global attackers.
@@ 3.4.4. Shuffler-specific attacks
On the other hand, TEE-based solutions can be sensitive to side-channel attacks (i.e., attacks exploiting leakage from the protocol implementation, but not the protocol itself) in which the memory access patterns help de-anonymize a shuffle of some constructions (Sasy et al., 2022).
@@ 4.1. Methodology
We adopted two approaches to identify relevant works able to implement a secure shuffler. First, we performed a survey on Google Scholar using different combinations of keyword including “secure shuffle,” “anonymous broadcaster,” “verifiable shuffler,” “mix network,” “anonymous communications,” “shuffle differential privacy,” and “secret shuffle.” This first step provided a preliminary list of secure shufflers. Second, we used Google Scholar to find additional secure shufflers in papers that cited these constructions. This second step enables to identify follow-up works that may propose more unconventional solutions to securely shuffle. For example, Bell et al. (Bell et al., 2020) introduced a secure-aggregation-based shuffler, while the main focus of their paper is secure aggregation.
@@
Fixed-route (Chaum, 1981; Danezis et al., 2003; Freedman and Morris, 2002) $O(1)$ $O(k)$ $O(n)$ $O(n)$ $\geq 1$
Free-route (Chaum, 1981; Dingledine et al., 2004; Chen et al., 2015) $O(1)$ $O(k)$ $O(nk/M)$ $O(nk/M)$ $\gg 1$
Advanced (Chaum et al., 2017; Piotrowska et al., 2017) $O(1)$ $O(k)$ $O(nk/M)$ $O(nk/M)$ $\gg 1$
Trellis (Langowski et al., 2023) $O(1)$ $O(k)$ $O(nk/M)$ $O(nk/M)$ $\gg 1$
TORaaS (Hohenberger et al., 2014) $O(1)$ $O(k)$ $O(n)$ $O(n)$ 1
@@
Classic (Furukawa and Sako, 2001; Neff, 2001; Bayer and Groth, 2012) $O(M)$ $O(1)$ $O(M\sqrt{n}+Mn)$ $O(n\log n+Mn+M\sqrt{n})$ $>1$
Riffle (Kwon et al., 2016) $O(1)$ $O(M)$ $O(n)$ $O(n)$ $>2$
Atom (Kwon et al., 2017) ($k$: group size, $\alpha=\frac{k}{M}$) $O(k)$ $O(1)$ $O(n\cdot\alpha+k\sqrt{n\cdot\alpha})$ $O(\alpha\log\alpha)$ $\gg 1$
Lattice-based (Aranha et al., 2023) $O(M)$ $O(1)$ $O(Mn)$ $O(Mn)$ $>1$
@@
Dissent (Corrigan-Gibbs and Ford, 2010) $O(Mn)$ $O(Mn)$ $O(n^{2}+nM)$ $O(Mn)$ $>1$
Riposte (Corrigan-Gibbs et al., 2015) $O(\sqrt{n})$ $O(\sqrt{n})$ $O(n\sqrt{n})$ $O(n^{2})$ $3$
Spectrum (Newman et al., 2022) $O(\log n)$ $O(\log n)$ $O(n\log n)$ $O(n^{2})$ 2
Sabre (Vadapalli et al., 2022) $O(\log n)$ $O(\log n)$ $O(n\log n)$ $O(n^{2})$ 3
@@
AsynchroMix (Lu et al., 2019b) $O(M)$ $O(M)$ $O(M\cdot n\cdot log^{2}n)$ $O(M\cdot n\cdot log^{2}n)$ $\geq 4$
PowerMix (Lu et al., 2019b) $O(M)$ $O(M)$ $O(n^{2})$ $O(n^{3})$ $\geq 4$
Blinder (Abraham et al., 2020) $O(M\sqrt{n})$ $O(M\sqrt{n})$ $O(n\sqrt{n})$ $O(n^{2})$ $\geq 5$
Clarion (Eskandarian and Boneh, 2022) $O(M)$ $O(M)$ $O(Mn)$ $O(Mn)$ $\geq 3$
RPM (Lu and Kate, 2023) $O(M)$ $O(M)$ $O(Mn\sqrt{n})$ $O(Mn^{2})$ $\geq 3$
Ruffle (A et al., 2023) $O(M)$ $O(M)$ $O(Mn)$ $O(Mn)$ $=3$
@@
Prochlo (Bittau et al., 2017) $O(1)$ $O(1)$ $O(n)$ $O(n)$ 1
ORShuffle (Sasy et al., 2022) $O(1)$ $O(1)$ $O(n)$ $O(n\log^{2}n)$ 1
Butterfly Shuffle (Gu et al., 2023) $O(1)$ $O(1)$ $O(n)$ $O(n\log n)$ 1
DBucket (Ngai et al., 2024) $O(1)$ $O(1)$ $O(n)$ $O(n\log n)$ 1
@@
Rumor-based (Fanti et al., 2017) ($\alpha=$ node degree) $O(\alpha\cdot n)$ $O(\alpha\cdot n)$ $O(\alpha\cdot n)$ $O(\alpha\cdot n)$ $\geq 1$
NIDAR (2 layers) (Bünz et al., 2021) $O(1)$ $O(1)$ $O(n)$ $O(n^{1+1/2})$ 1
Secure-Agg.-based (Bell et al., 2020) $O(log^{2}n+n\log n)$ $O(n\log^{2}n)$ $O(n^{2}\log n)$ $O(n^{2}\cdot\log^{2}n)$ 1
@@
Fixed-route (Chaum, 1981; Danezis et al., 2003; Freedman and Morris, 2002) $>1$ $M-1$ malicious servers
Free-route (Chaum, 1981; Dingledine et al., 2004; Chen et al., 2015) $\gg 1$ $k-1$ malicious servers per path De-anonymization atk.
Advanced (Chaum et al., 2017; Piotrowska et al., 2017) $\gg 1$ $k-1$ malicious servers per path
Trellis (Langowski et al., 2023) $\gg 1$ $f$of malicious servers Trusted setup & Weaker anonymity distribution
TORaaS (Hohenberger et al., 2014) 1 $k-1$ malicious servers per path De-anonymization atk.
@@
Classic (Furukawa and Sako, 2001; Neff, 2001; Bayer and Groth, 2012) $>1$ $M-1$ malicious servers Trusted setup
Riffle (Kwon et al., 2016) $>2$ $M-2$ malicious servers Frequent setup w/ (Bayer and Groth, 2012)
Atom (Kwon et al., 2017) ($k$: group size, $\alpha=\frac{k}{M}$) $\gg 1$ $k-1$ malicious serv. per group Possibly a weaker anonymity distribution
Lattice-based (Aranha et al., 2023) $>1$ $M-1$ malicious servers Trusted setup
@@
Dissent (Corrigan-Gibbs and Ford, 2010) $>1$ $M-1$ malicious servers Setup w/ (Bayer and Groth, 2012)
Riposte (Corrigan-Gibbs et al., 2015) $3$ 1 malicious out of 3 servers
Spectrum (Newman et al., 2022) $2$ $M-1$ malicious servers Setup uses another secure shuffler
Sabre (Newman et al., 2022) $3$ $M-1$ malicious servers
@@
AsynchroMix (Lu et al., 2019b) $\geq 4$ $t<M/3$ malicious
PowerMix (Lu et al., 2019b) $\geq 4$ $t<M/3$ malicious
Blinder (Abraham et al., 2020) $\geq 5$ $t<M/4$ malicious
Clarion (Eskandarian and Boneh, 2022) $\geq 3$ $M-1$ malicious
RPM (Lu and Kate, 2023) $\geq 3$ $t<M/2$ malicious
Ruffle (A et al., 2023) $=3$ $t<M/2$ malicious
@@
Prochlo (Bittau et al., 2017) $1$ Malicious server Trusted hardware & security issues (Sasy et al., 2022)
ORShuffle (Sasy et al., 2022) $1$ Malicious server Trusted hardware
Butterfly Shuffle (Gu et al., 2023) $1$ Malicious server Trusted hardware
DBucket (Ngai et al., 2024) $1$ Malicious server Trusted hardware
@@
Rumor-based (Fanti et al., 2017) ($\alpha=$ node degree) $\geq 1$ $t_{\alpha}$ malicious agents out of $N+M$ De-anonymization attacks
NIDAR (2 layers) (Bünz et al., 2021) $1$ Malicious server Trusted setup & Weaker anonymity distribution
Sec.-Agg.-based (Bell et al., 2020) $1$ Malicious server
@@ 4.2. Mix networks (mix nets)
Mix networks are a well-known anonymization technique thanks to Tor (Dingledine et al., 2004) (a popular mix network used notably by activists and journalists for anonymous Web browsing). The main idea of mix networks (Chaum, 1981) is to make a packet transit through several relay nodes before reaching its destination. The packets are “onion-encrypted”: one layer of encryption is added for each relay in the path. Upon receiving the packet, each relay node removes one layer of encryption and forwards the decrypted payload to the next relay. This mixing process prevents the relay nodes from linking the sender to the receiver. For secure shuffling, all the messages are sent to the same agent (i.e., the analyzer or a public bulletin).
@@
Remark about semi-trusted shuffler. Some papers, such as (Humphries et al., 2022), use a semi-trusted shuffler that acts as a proxy. A mix network being a proxy chain, the semi-trusted shuffler is a 1-node mix network with a non-colluding receiver.
(a) Fixed-route
[…]
[Schema of a fixed-route mix net]Schema of a fixed-route mix net: three data owners send their encrypted message to a chain of servers, the servers successively shuffles the ciphertexts and transmit the output to an analyzer.
(b) Free-route
[…]
[Schema of a free-route mix net]Schema of a free-route mix net: three data owners send their encrypted message through a network of server, each message takes a different pass and finishes on an analyzer.

### v2

@@ 3.3.2. Correctness
Next to anonymity, a good shuffler should also satisfy correctness. This property covers various security definitions present in existing security papers. On the one hand, secure shufflers based on MPC (e.g., RPM (lu_rpm_2023)) prove a standard security property called “correctness” (i.e., the MPC protocol outputs the expected output). On the other hand, recent anonymous communication systems (e.g., Blinder (abraham_blinder_2020)) often guarantee censorship resistance. This second security guarantee is essential, but it does not prevent all the threats covered by the “correctness.” Taking inspiration from MPC, we propose the following explicit definition for shuffling correctness:
@@ 3.3.3. Sender disruption resistance
Several ML works studied the resilience of distributed training protocols to malicious behaviors (chen_distributed_2017; alistarh_byzantine_2018). These works conceive protocols that succeed even in the presence of malicious or faulty agents. Protocol disruption is also possible in secure shuffling and is critical when thousands or millions of senders are involved. Controlling the honesty of millions of senders is too complex, so a form of disruption resistance is necessary. We propose the following definition regarding the resistance to malicious senders:
@@
This property provides tolerance to drop-outs (i.e., the disconnections of senders before the protocol completion), which drew much attention in the secure aggregation literature (bell_secure_2020; sabater_accurate_2021; bell_acorn_2022). Secure aggregation protocols such as (bell_secure_2020; bell_acorn_2022) only tolerate a fixed number of drop-outs, while sender disruption-resistant shufflers tolerate an unbounded number of drop-outs. Sender disruption resistance is stronger than drop-out tolerance since it also covers the reception of ill-formed data.
@@
Finally, Definition 3.7 specifies that the disruption can only be caused by missing or ill-formed messages to exclude Distributed Denial-of-Service (DDoS) attacks (i.e., flooding a server with a massive amount of requests to make it unavailable). These attacks (or any other network-level attack) are not specific to shufflers and can be mitigated using network security techniques (mirkovic_taxonomy_2004).
@@ 3.3.4. Server disruption resistance
This property is called “robustness” in MPC protocols (abraham_blinder_2020). One may see it as an extra property because the number of servers is small enough to maintain a blocklist of untrustworthy servers. As for sender disruption resistance, server disruption resistance brings resilience to drop-outs.
@@ 3.4.1. Intersection attacks
These attacks (wolinsky_hang_2013; danezis_statistical_2005) rely on the statistical analysis that can be made from several shuffle outputs to de-anonymize messages. Indeed, if a data owner participates in $k$ shuffles, computing the intersection of the round results might leak some information about this data owner. These attacks against shufflers are specific to the use case and the data being shuffled. For example, the intersection attack is useless if the shuffled messages are secret shares, as in Ishai et al. (ishai_cryptography_2006) protocol, but the intersection may de-anonymize the messages if we shuffle cleartext data.
Wolinsky et al. (wolinsky_hang_2013) propose a generic Buddy system compliant with any anonymization system to mitigate these attacks. The general idea is to ensure a large set of possible senders for a given message stream, so the intersection of multiple shuffles cannot lead to a de-anonymization. No result is available regarding the scalability of this generic mitigation technique for applications to privacy-preserving computations.
@@ 3.4.2. Trickle and flooding attacks
Serjantov et al. (goos_trickle_2003) describe two general attacks (grouped under the term “blending attacks”): trickle attacks and flooding attacks.
@@
The trickle attack is only possible on a system insecure against censorship (i.e., with no correctness guarantee as defined in Section 3). The flooding attack is only possible if a malicious server colludes with data owners and can be mitigated using an authentication scheme, so a single malicious user cannot flood a batch alone. Some secure shufflers (e.g., mix nets) must use anonymous authentication systems (chase_algebraic_2014) to mitigate these attacks.
@@ 3.4.3. Traffic analysis attacks
These attacks (murdoch_low-cost_2005; sun_raptor_2015; arp_torben_2015) have been extensively documented for mix networks: a global adversary (e.g., a state agency) observes the packet transmission between the agents to trace the path of the messages. From the beginning, anonymization networks such as Tor were presented as possibly insecure against global attackers.
@@ 3.4.4. Shuffler-specific attacks
On the other hand, TEE-based solutions can be sensitive to side-channel attacks (i.e., attacks exploiting leakage from the protocol implementation, but not the protocol itself) in which the memory access patterns help de-anonymize a shuffle of some constructions (sasy_fast_2022).
@@ 4.1. Methodology
We adopted two approaches to identify relevant works able to implement a secure shuffler. First, we performed a survey on Google Scholar using different combinations of keyword including “secure shuffle,” “anonymous broadcaster,” “verifiable shuffler,” “mix network,” “anonymous communications,” “shuffle differential privacy,” and “secret shuffle.” This first step provided a preliminary list of secure shufflers. Second, we used Google Scholar to find additional secure shufflers in papers that cited these constructions. This second step enables to identify follow-up works that may propose more unconventional solutions to securely shuffle. For example, Bell et al. (bell_secure_2020) introduced a secure-aggregation-based shuffler, while the main focus of their paper is secure aggregation.
@@
Fixed-route (chaum1981untraceable; danezis_mixminion_2003; freedman_tarzan_2002) $O(1)$ $O(k)$ $O(n)$ $O(n)$ $\geq 1$
Free-route (chaum1981untraceable; dingledine2004tor; chen_hornet_2015) $O(1)$ $O(k)$ $O(nk/M)$ $O(nk/M)$ $\gg 1$
Advanced (chaum_cmix_2017; piotrowska_loopix_2017) $O(1)$ $O(k)$ $O(nk/M)$ $O(nk/M)$ $\gg 1$
Trellis (langowski_trellis_2023) $O(1)$ $O(k)$ $O(nk/M)$ $O(nk/M)$ $\gg 1$
TORaaS (hohenberger_anonize_2014) $O(1)$ $O(k)$ $O(n)$ $O(n)$ 1
@@
Classic (furukawa_efficient_2001; neff_verifiable_2001; bayer_efficient_2012) $O(M)$ $O(1)$ $O(M\sqrt{n}+Mn)$ $O(n\log n+Mn+M\sqrt{n})$ >1>1
Riffle (kwon_riffle_2016) $O(1)$ $O(M)$ $O(n)$ $O(n)$ >2>2
Atom (kwon_atom_2017) ($k$: group size, $\alpha=\frac{k}{M}$) $O(k)$ $O(1)$ $O(n\cdot\alpha+k\sqrt{n\cdot\alpha})$ $O(\alpha\log\alpha)$ $\gg 1$
Lattice-based (aranha_verifiable_2023) $O(M)$ $O(1)$ $O(Mn)$ $O(Mn)$ >1>1
@@
Dissent (corrigan-gibbs_dissent_2010) $O(Mn)$ $O(Mn)$ $O(n^{2}+nM)$ $O(Mn)$ >1>1
Riposte (corrigan-gibbs_riposte_2015) $O(\sqrt{n})$ $O(\sqrt{n})$ $O(n\sqrt{n})$ $O(n^{2})$ 33
Spectrum (newman_spectrum_2022) $O(\log n)$ $O(\log n)$ $O(n\log n)$ $O(n^{2})$ 2
Sabre (vadapalli_sabre_2022) $O(\log n)$ $O(\log n)$ $O(n\log n)$ $O(n^{2})$ 3
@@
AsynchroMix (lu_honeybadgermpc_2019) $O(M)$ $O(M)$ $O(M\cdot n\cdot log^{2}n)$ $O(M\cdot n\cdot log^{2}n)$ $\geq 4$
PowerMix (lu_honeybadgermpc_2019) $O(M)$ $O(M)$ $O(n^{2})$ $O(n^{3})$ $\geq 4$
Blinder (abraham_blinder_2020) $O(M\sqrt{n})$ $O(M\sqrt{n})$ $O(n\sqrt{n})$ $O(n^{2})$ $\geq 5$
Clarion (eskandarian_clarion_2022) $O(M)$ $O(M)$ $O(Mn)$ $O(Mn)$ $\geq 3$
RPM (lu_rpm_2023) $O(M)$ $O(M)$ $O(Mn\sqrt{n})$ $O(Mn^{2})$ $\geq 3$
Ruffle (a_ruffle_2023) $O(M)$ $O(M)$ $O(Mn)$ $O(Mn)$ =3=3
@@
Prochlo (bittau_prochlo_2017) $O(1)$ $O(1)$ $O(n)$ $O(n)$ 1
ORShuffle (sasy_fast_2022) $O(1)$ $O(1)$ $O(n)$ $O(n\log^{2}n)$ 1
Butterfly Shuffle (gu_efficient_2023) $O(1)$ $O(1)$ $O(n)$ $O(n\log n)$ 1
DBucket (ngai_distributed_2024) $O(1)$ $O(1)$ $O(n)$ $O(n\log n)$ 1
@@
Rumor-based (fanti_hiding_2017) ($\alpha=$ node degree) $O(\alpha\cdot n)$ $O(\alpha\cdot n)$ $O(\alpha\cdot n)$ $O(\alpha\cdot n)$ $\geq 1$
NIDAR (2 layers) (bunz_non-interactive_2021) $O(1)$ $O(1)$ $O(n)$ $O(n^{1+1/2})$ 1
Secure-Agg.-based (bell_secure_2020) $O(log^{2}n+n\log n)$ $O(n\log^{2}n)$ $O(n^{2}\log n)$ $O(n^{2}\cdot\log^{2}n)$ 1
@@
Fixed-route (chaum1981untraceable; danezis_mixminion_2003; freedman_tarzan_2002) $\geq 1$ M1M-1 malicious servers
Free-route (chaum1981untraceable; dingledine2004tor; chen_hornet_2015) $\gg 1$ k1k-1 malicious servers per path De-anonymization atk.
Advanced (chaum_cmix_2017; piotrowska_loopix_2017) $\gg 1$ k1k-1 malicious servers per path
Trellis (langowski_trellis_2023) $\gg 1$ ff% of malicious servers Trusted setup & Weaker anonymity distribution
TORaaS (hohenberger_anonize_2014) 1 k1k-1 malicious servers per path De-anonymization atk.
@@
Classic (furukawa_efficient_2001; neff_verifiable_2001; bayer_efficient_2012) >1>1 M1M-1 malicious servers Trusted setup
Riffle (kwon_riffle_2016) >2>2 M2M-2 malicious servers Frequent setup w/ (bayer_efficient_2012)
Atom (kwon_atom_2017) ($k$: group size, $\alpha=\frac{k}{M}$) $\gg 1$ k1k-1 malicious serv. per group Possibly a weaker anonymity distribution
Lattice-based (aranha_verifiable_2023) >1>1 M1M-1 malicious servers Trusted setup
@@
Dissent (corrigan-gibbs_dissent_2010) >1>1 M1M-1 malicious servers Setup w/ (bayer_efficient_2012)
Riposte (corrigan-gibbs_riposte_2015) 33 1 malicious out of 3 servers
Spectrum (newman_spectrum_2022) 22 M1M-1 malicious servers Setup uses another secure shuffler
Sabre (newman_spectrum_2022) 33 M1M-1 malicious servers
@@
AsynchroMix (lu_honeybadgermpc_2019) $\geq 4$ $t<M/3$ malicious
PowerMix (lu_honeybadgermpc_2019) $\geq 4$ $t<M/3$ malicious
Blinder (abraham_blinder_2020) $\geq 5$ $t<M/4$ malicious
Clarion (eskandarian_clarion_2022) $\geq 3$ M1M-1 malicious
RPM (lu_rpm_2023) $\geq 3$ $t<M/2$ malicious
Ruffle (a_ruffle_2023) =3=3 $t<M/2$ malicious
@@
Prochlo (bittau_prochlo_2017) 11 Malicious server Trusted hardware & security issues (sasy_fast_2022)
ORShuffle (sasy_fast_2022) 11 Malicious server Trusted hardware
Butterfly Shuffle (gu_efficient_2023) 11 Malicious server Trusted hardware
DBucket (ngai_distributed_2024) 11 Malicious server Trusted hardware
@@
Rumor-based (fanti_hiding_2017) ($\alpha=$ node degree) $\geq 1$ $t_{\alpha}$ malicious agents out of $N+M$ De-anonymization attacks
NIDAR (2 layers) (bunz_non-interactive_2021) 11 Malicious server Trusted setup & Weaker anonymity distribution
Sec.-Agg.-based (bell_secure_2020) 11 Malicious server
@@ 4.2. Mix networks (mix nets)
Mix networks are a well-known anonymization technique thanks to Tor (dingledine2004tor) (a popular mix network used notably by activists and journalists for anonymous Web browsing). The main idea of mix networks (chaum1981untraceable) is to make a packet transit through several relay nodes before reaching its destination. The packets are “onion-encrypted”: one layer of encryption is added for each relay in the path. Upon receiving the packet, each relay node removes one layer of encryption and forwards the decrypted payload to the next relay. This mixing process prevents the relay nodes from linking the sender to the receiver. For secure shuffling, all the messages are sent to the same agent (i.e., the analyzer or a public bulletin).
@@
Remark about semi-trusted shuffler. Some papers, such as (humphries_selective_2022), use a semi-trusted shuffler that acts as a proxy. A mix network being a proxy chain, the semi-trusted shuffler is a 1-node mix network with a non-colluding receiver.
(a) Fixed-route: Schema of a fixed-route mix net
(b) Free-route: Schema of a free-route mix net